The perimeter of a rectangular plot of land is 250 meters, what is the area of the plot, in square meters?
Answer by checkley75(3666)

IF IT IS A SQUARE THEN: 4S=250 S=250/4 S=62.5 FOR EACH SIDE.
THEN THE AREA=62.5^62.5=3,906.25 SQ. METERS.
ANY OTHER RECTANGLE LIKE A 50 BY 75 METER RECTANGLE WE GET
50*75=3,750 SQ. METERS.
-------------------------------------------------------------------
TO YOUR GENERAL QUAESTION -- THERE IS NO ONE ANSWER.